Let f(x) =
x-2
_____
8
and g(x) = 8x+2
find (f o g)(x).
Please show the work so I can see how the problem was worked out. Thanks

Answer by rothauserc(4718)

f(x) = (x-2) / 8
g(x) = 8x +2
(f o g)(x) = f(g(x))
(f o g)(x) = f(8x+2) = (8x+2-2) / 8
combine terms in numerator
(f o g)(x) = 8x / 8
divide numerator and denominator by 8
(f o g)(x) = x